How far will a truck travelling at 80km/h travel in 2h 45 min?
Answer by lenny460(1073)

How far will a truck travelling at 80km/h travel in 2h 45 min?
In 1 minute the truck will travel:
80 / 60 = 1.33 km
In 45 minutes the truck will travel:
45 * 1.33 = 59.85 km
In 2 hours the truck will travel:
80 * 2 = 160 km

Therefore:
160 + 59.85 = 219.85 km

Answer:
219.85 km



Lennox Obuong
Algebra Tutor
Nairobi, Kenya